### v1

# Topology Sculptor, Shape Refiner: Discrete Diffusion Model for High-Fidelity 3D Meshes Generation

Abstract
In this paper, we introduce Topology Sculptor, Shape Refiner (TSSR), a novel method for generating high-quality, artist-style 3D meshes based on Discrete Diffusion Models (DDMs). Our primary motivation for TSSR is to achieve highly accurate token prediction while enabling parallel generation, a significant advantage over sequential autoregressive methods. By allowing TSSR to ”see” all mesh tokens concurrently, we unlock a new level of efficiency and control. We leverage this parallel generation capability through three key innovations: 1) Decoupled Training and Hybrid Inference, which distinctly separates the DDM-based generation into a topology sculpting stage and a subsequent shape refinement stage. This strategic decoupling enables TSSR to effectively capture both intricate local topology and overarching global shape. 2) An Improved Hourglass Architecture, featuring bidirectional attention enriched by face-vertex-sequence level Rotational Positional Embeddings (RoPE), thereby capturing richer contextual information across the mesh structure. 3) A novel Connection Loss, which acts as a topological constraint to further enhance the realism and fidelity of the generated meshes. Extensive experiments on complex datasets demonstrate that TSSR generates high-quality 3D artist-style meshes, capable of achieving up to 10,000 faces at a remarkable spatial resolution of $1024^{3}$. The code will be released at: https://github.com/psky1111/Tencent-TSSR.
1 Introduction
High-fidelity 3D mesh generation [28, 5] has garnered significant attention, focusing on generating the desired shape and high-quality topology. It produces detailed, topologically sound geometric representations for a wide range of applications, including virtual reality, digital content creation, and robotic simulation.
Mainstream 3D mesh generation methods can be broadly categorized into two approaches: 1) Shape-then-mesh generation and 2) direct mesh generation. Shape-then-mesh generations [28] start with representing a shape via continuous volumetric fields, such as Neural Radiance Fields (NeRF) [17] and Signed Distance Fields (SDF) [10]. Then, a diffusion model is used to generate such a representation and convert it to real meshes via Marching Cubes [15]. Benefits from parallel token generation in diffusion models, but shape-then-mesh generation is efficient but suffers from inevitable denegation after the conversion process. Direct mesh generation methods aim to predict mesh vertices and faces directly. Mainstream methods, such as Meshtron [8] and DeepMesh [27], employ the Auto-Regressive (AR) approach, which first quantizes the mesh coordinates into a discrete space (V discrete bins for each coordinate, where V is also referred to as the spatial resolution in this context). Then it implements the specific tokenizer method to generate the mesh token-by-token. Benefits from the discrete space, ARs can generate high-quality meshes, but struggle with efficiency.
Discrete Diffusion Models (DDMs) serve as a remedy to enable parallel generation in a discrete space, such as SeedDiffusion [21] and LLaDA [18] in large language models (LLMs). DDMs follow a similar process to the diffusion models, which contain: 1) the forward process, adding noise to the discrete token sequence, and 2) the reverse process, denoising the noise-corrupted token sequence. However, token-parallel generation inherently breaks the causal sequence (the generated order), leading to less accurate token prediction when encountering long, high-resolution sequences. In 3D mesh generation, this problem is vastly magnified, as the causal sequence can directly ensure valid topology and precise token prediction [8, 5]. A natural question thus arises: ”How to maintain accurate token prediction in long and high spatial sequences when using parallel generation (which breaks causal order)?”
In this paper, we propose the Topology Sculptor, Shape Refiner (TSSR), a novel paradigm explicitly designed for direct mesh generation built upon DDMs. TSSR addresses the challenge of maintaining accurate token prediction in parallel generation for meshes. TSSR leverages DDMs’ inherent ability for parallel token generation, which enables the model to view all tokens simultaneously and utilize bidirectional attention to capture comprehensive global shape and intricate topological information. TSSR realizes accurate parallel mesh generation through three integrated core contributions:
1) A unique decoupling training and hybrid inference paradigm that strategically divides mesh generation into specialized sub-tasks for shape (shape refiner) and topology (topology sculptor). This strategy acknowledges that while parallel generation yields rich, comprehensive information (encompassing global shape and intricate topology), effectively processing this intertwined complexity for accurate prediction is challenging. Decoupling training enables the separation of these aspects into distinct sub-tasks, thereby reducing the learning burden on the model and allowing each sub-task to focus predominantly on either shape or topology. Then, the hybrid inference process unifies these sub-tasks into a complete, iterative refinement loop, where mutually enhancing shape and topology prediction is achieved, thereby improving token prediction accuracy. 2) An improved Hourglass architecture designed to process and leverage multi-level mesh information effectively. This architecture integrates Multi-level RoPE and an enhanced transformer block to capture crucial hierarchical geometric context, supporting highly accurate token prediction across long and spatially complex sequences. 3) A novel connection loss that directly introduces topology priors during training. This loss enforces hard topology constraints by ensuring consistency for tokens sharing the same vertex, thereby significantly enhancing the model’s ability to maintain topological integrity. Collectively, these three components ensure that TSSR fully utilizes the rich information available in parallel token generation. This strategic integration significantly improves token prediction accuracy, thereby addressing the loss of sequential causal dependencies and rendering parallel mesh generation highly effective for artist-style meshes. Experimental results on various open datasets demonstrate TSSR’s successful generation of artist-style meshes via parallel token generation at high resolutions (i.e., $1024^{3}$) up to 10K faces.
To summarize, our key contributions are as follows:
  • We propose the Topology Sculptor, Shape Refiner, a novel DDM-based paradigm specifically tailored for artist-style 3D mesh generation.
  • By carefully designing, TSSR enables full leverage of the rich information after seeing all tokens of models in token parallel generation.
  • We demonstrate that TSSR achieves state-of-the-art performance in generating high-quality 3D artist-style meshes with up to 10,000 faces at a spatial resolution of $1024^{3}$, showcasing superior global coherence and local precision.
2 Related Work
Shape-then-mesh generation. These works focus on first generating the shape via a continuous volumetric field, where SDF is the most common choice. Then, SDF can be converted to the mesh via the Marching Cube [15]. For example, TRELLIS [25] and Hunyuan 2.5 [11] first leverage the VAE to reconstruct the mesh via SDF. Then, they leverage the diffusion model to generate the VAE latent. The advantage of this is that the explicitly 3D information can be introduced into the VAE. Diffusion itself can be more robust to the 3D shape. The difference is that TRELLIS introduces the explicit shape constraint into VAE training. Hunyuan proposed the improved version to extract GT SDF. Besides enhancing VAE, Wonder3D [14] introduces a cross-domain diffusion model framework to further enhance the diffusion process.
Direct mesh generation. Direct mesh generation aims at generating a mesh without conversion, such as SDF to mesh. The mainstream way is the ARs method. For example, MeshGPT [20] pioneers AR-based mesh synthesis through sequence-based modeling. MeshAnything [4] and MeshAnything V2 [5] proposed an improved tokenizer through adjacent mesh tokenization. MeshXL [3] introduced the Neural Coordinate Field, which fuses explicit coordinate representation with implicit neural embeddings to enhance the AR. EdgeRunner [22] proposed compressing variable-length meshes into fixed-size latent vectors to reduce the sequence length, which improves the efficiency of ARs. Meshtron [8] proposed the Hourglass architecture to enhance mesh generation quality and efficiency further. MeshMosaic [26] proposed the part-splitting-based method to allow parallel generation of the part.
Except for the ARs, there is some exploration for token parallel generation via diffusion. For example, Meshcraft [9] and Polydiff [1] proposed directly mapping the mesh into a continuous latent space via VAE and then implementing diffusion models on the latent space. However, the compression loss of VAE will inevitably lead to the degradation again.
Discrete Diffusion Models. DDMs [7, 19] are another type of diffusion model compared to the continuous state-based diffusion models [13]. Previous works have begun to implement DDMs on various tasks, such as LLM [21] and image generation [23], demonstrating the potential of DDMs. In this work, we propose TSSR, a token parallel generation paradigm based on the DDM. TSSR directly generates the artist-like method without any conversion, such as SDF and VAE, thereby maintaining the high quality.
3 Preliminaries
This section provides essential background on discrete mesh representation and the general framework of DDMs as applied to mesh generation.
Discrete representation of mesh. A 3D mesh $M$ is fundamentally represented by a set of $N$ faces, denoted as $\{\text{f}_{i}\}_{i=1}^{N}$. Each face $\text{f}_{i}$ is composed of $n$ vertices, $\{\text{v}_{i,j}\}_{j=1}^{n}$. In this paper, we specifically consider triangular meshes, thus setting $n=3$, where each face has three vertices. Each vertex $\text{v}_{i,j}$ is defined by its three-dimensional Cartesian coordinates: $\{v_{i,j}x,v_{i,j}y,v_{i,j}z\}$. To enable processing by discrete generative models, vertex coordinates are first discretized into a fixed number of bins, defining the spatial resolution. Consistent with the high-fidelity demands of artist-style mesh generation, we adopt a spatial resolution of $1024^{3}$, meaning coordinates are quantized into $V=1024$ discrete bins along each axis. For discrete diffusion models on meshes, these discretized coordinates are transformed into a sequence of tokens. We utilize the Meshtron tokenizer [8], which flattens all vertex coordinates of a mesh into a one-dimensional sequence. Specifically, a tokenized mesh $\hat{x}_{1}$ is represented as:
$\hat{x}_{1}=\{v_{1,1}x,v_{1,1}y,v_{1,1}z,\ldots,v_{N,n}x,v_{N,n}y,v_{N,n}z\}.$ (1)
The sequence length of $\hat{x}_{1}$ is $9N$, as each of the $N$ triangular faces contributes 9 coordinate tokens (3 vertices $\times$ 3 coordinates per vertex) [8]. The effective codebook size for these tokens is $B=V+k$, where $V=1024$ represents the quantized coordinate values and $k$ denotes the number of any additional special tokens (e.g., start-of-sequence, end-of-sequence, padding) used in the tokenization scheme.
Discrete Diffusion Models (DDMs) Discrete Diffusion Models (DDMs) are a class of generative paradigms that learn to transform pure noise into target data via discrete space. This is achieved by defining a forward diffusion process that gradually corrupts clean data ($\hat{x}_{1}$) into noise ($x_{0}$), and then learning to reverse this process to generate data. Unlike continuous diffusion models that often define a velocity field in a continuous space, DDMs are characterized by a model that predicts the clean target state, $p_{1|t}(x_{t},t,c,\theta)$, from a noisy state $x_{t}$ at time $t\in[0,1]$ (where $t=0$ represents initial noise and $t=1$ represents the clean state).
The core of DDM inference lies in defining an iterative update rule that transitions from a current noisy state $x_{t}$ to a slightly cleaner state $x_{t+\Delta t}$ (or $x_{t-\Delta t}$), leveraging the predicted clean state $p_{1|t}$. Such a transition could be achieved by two views:Probability Space (e.g., Discrete Flow Matching [7]). $p_{1|t}$ is used to define a continuous-time velocity field in a probability space, which then guides the transition from $x_{t}$ to $x_{t+\Delta t}$. This often involves complex mathematical derivations for the update rule.
Token Space Updates (e.g., as in SeedDiffusion [21] and TSSR ): In this approach, $p_{1|t}(x_{t},t,c,\theta)$ is directly interpreted as a categorical distribution over tokens. The update rule for $x_{t}\to x_{t+\Delta t}$ is then constructed by sampling from this $p_{1|t}$ prediction to decide which clean tokens to incorporate, or which tokens to mask/randomize, to form the next state $x_{t+\Delta t}$. This method operates directly in the discrete token space, making it more intuitive.
In DDMs, two primary types of noise are commonly used to define the corruption process in the forward pass:
  1. 1.
    Mask-based Noise: Corrupts tokens by replacing ground truth values with a special mask, typically denoted as $\text{MASK}]$. This formulation inherently supports inpainting tasks, as the model learns to fill in missing information based on surrounding context.
  2. 2.
    Uniform-based Noise: Corrupts tokens by replacing ground truth values with random tokens sampled uniformly from the entire valid codebook. This approach encourages the model to learn global structural recovery, as it must reconstruct the data from entirely random inputs.
The forward process of a DDM defines the distribution of a noisy state $x_{t}$ at time $t\in[0,1]$, given the clean data $\hat{x}_{1}$ and the related noise data $x_{0}$:
$p(x_{t}|\hat{x}_{1},x_{0})=\kappa_{t}\delta_{x_{0}}(x_{0})+(1-\kappa_{t})\delta_{\hat{x}_{1}}(\hat{x}_{1}),$ (2)
where $\kappa(t)$ is a time-dependent noise scheduler that controls how many tokens in $\hat{x}_{1}$ will be replaced by noise tokens $x_{0}$. $\delta_{x}(y)$ is a Dirac delta function (or its discrete equivalent, a one-hot distribution) indicating that $x=y$.
The inference process in DDMs aims to generate samples from $p(\hat{x}_{1})$ by starting from a noisy state (e.g., pure mask tokens or random tokens) and iteratively refining it from $t=0$ (noise) to $t=1$ (clean) or vice versa. The model predicts the clean target state, $p_{1|t}(x_{t},t,c,\theta)$, and this prediction is then used by a custom update rule to transition to the next state in the refinement sequence.
The training process for DDMs optimizes the model parameters $\theta$ to predict the final clean state. This is commonly achieved by minimizing a cross-entropy loss between the model’s prediction of the clean state and the actual clean state $\hat{x}_{1}$:
$\mathcal{L}_{DDM}=\text{CE}(p_{1|t}(x_{t},t,c,\theta),\hat{x}_{1}),$ (3)
where $x_{t}$ is sampled from the forward process given $\hat{x}_{1}$.
4 Method
In this section, we demonstrate how TSSR effectively leverages the global shape and intricate topology available in parallel token generation. This is achieved through three integrated core contributions: 1) a unique decoupling training and hybrid inference, 2) an improved Hourglass architecture, and 3) the proposed connection loss.
Decoupling via inherent noise biases. The TSSR paradigm is motivated by the distinct learning biases inherent in the two primary DDMs noise types, which we leverage to decouple the complex task of mesh generation.
  • Mask-based noise for topology sculpting: As an infilling task, denoising mask-corrupted data naturally drives the model to focus on local structural completion, such as maintaining connectivity and filling small holes. This inherent behavior biases the model to act as a powerful topology sculptor.
  • Uniform-based noise for shape refining: Denoising from a ”totally broken shape” of random valid tokens forces the model to recover the overall global form and proportions, as local structural priors are destroyed. This naturally biases the model to function as a powerful shape refiner.
Figure 1: An illustration of different types of noise for meshes. Then [MASK] is treated as a token removal. As a result, mask-based noise will force the model to focus on local structural completion. Uniform noise will generate a ”hedgehog”-like mesh, leading the model to focus on the shape.
We visualize the corrupted mesh via two noises, as shown in Fig. 1, to further illustrate the learning biases. Leveraging this insight, TSSR develops a hybrid process that combines these two specialized roles to achieve both accurate topology and coherent shape.
Hybrid inference and decoupling training paradigm. Firstly, we demonstrate the hybrid inference process to clearly show how TSSR transfers the state from $x_{t}$ to $x_{t+\Delta t}$. We assume there are two pre-trained models, $\theta_{mask}$ and $\theta_{uniform}$, which are responses to the topology sculptor and the shape refiner, respectively. Then, the inference process involves: 1) the initial states, which decide what type of noise in $x_{t}$ and $x_{t+\Delta t}$, and 2) the iterative update rule from $x_{t}$ to $x_{t+\Delta t}$. The inference process starts with the pure mask token-based sequence, i.e., $x_{0}$ is composed of [MASK]. This is the trick from the theoretical gargantuan to enable customization of the overall inference process and to ignore the rates of GT and noise tokens during training [7]. Therefore, $x_{t}$ and $x_{t+\Delta t}$ are the masked-noise-corrupted versions.
The iterative rule involves cycling through the topology sculptor $\theta_{mask}$ and $\theta_{uniform}$, as well as a classifier $\phi$, to decide which tokens to keep. We define $x_{t,mask}$ and $x_{t,uniform}$ to represent the mask-based noise corrupted and uniform-based noise corrupted $x_{t}$, respectively. Then, $x_{t}$, $t<1$ , always have mask noise and thus $x_{t,massk}=x_{t}$. We do the mask-based denoising process to calculate $p_{1|t}(x_{t,mask},t,c,\theta_{mask})$ and get its related clean state prediction $x_{1|t,mask}$. $x_{1|t,mask}$ will inevitably have some errors in token results due to the bias of $\theta_{mask}$, but keep some local intricate topology based on our motivation. Thus, it can be naturally regarded as adding a special type of uniform-based noise caused by model bias. This allows TSSR to use the shape refiner to correct these error tokens again and recover the shape, where we have $x_{t,uniform}=x_{1|t,mask}$. We then do uniform-based denoising process to calculate $p_{1|t}(x_{t,uniform},t,c,\theta_{uniform})$ and get its related clean state prediction $x_{1|t,uniform}$.
Now, the final step is to decide how many tokens should be kept and transfer the rest to [MASK] to get $x_{t+\Delta t}$. We leverage $\phi$ to calculate:
$\text{Softmax}(p_{\phi}(x_{1|t,uniform}))<\sigma,$ (4)
where $\text{Softmax}(\ast)$ is the softmax operator, $p_{\phi}(x_{1|t,uniform})$ is the logits about whether the tokens are correct via classifier $\phi$. Eq. 4 defines the confidence about the $x_{1|t,uniform}$. For each token in $x_{1|t,uniform}$, the confidence below than $\sigma$ will be replaced by [MASK]. In this way, we can $x_{t+\Delta t}$ after re-masking by Eq. 4 and finish one iteration via our hybrid inference.
Then, we demonstrate our decoupling training. Hybrid inference are related to three models $\theta_{mask}$, $\theta_{uniform}$ and $\phi$. For $\theta_{mask}$, it follows the standard mask noise-based DDMs. We generate $x_{t,mask}$ via Eq. 2. Then, the $\theta_{mask}$ is to predicted the clean state $p_{1|t}(x_{t,mask},t,c,\theta_{mask})$, where the loss can be formulated via Eq. 3 as:
$\mathcal{L}_{mask}=\text{CE}(p_{1|t}(x_{t},t,c,\theta_{mask}),\hat{x}_{1}).$ (5)
Then, for $\theta_{uniform}$, we follow the hybrid inference to calculate $x_{t,uniform}$ via $x_{t,uniform}\sim p_{1|t}(x_{t,mask},t,c,\theta_{mask})$ and the loss can be formulated:
$\mathcal{L}_{uniform}=\text{CE}(p_{1|t}(x_{t},t,c,\theta_{uniform}),\hat{x}_{1}).$ (6)
Similarly, we directly calculate $p_{\phi}(x_{1|t,uniform})$ via $x_{1|t,uniform}\sim p_{1|t}(x_{t},t,c,\theta_{uniform})$ and build a binary classification task via:
$\mathcal{L}_{\phi}=\text{BE}(p_{\phi}(x_{1|t}),\hat{x}_{1}),$ (7)
where $\text{BE}(\ast)$ is the binary classification loss.
In this way, the overall loss can be defined as follows:
$\mathcal{L}=\mathcal{L}_{mask}+\mathcal{L}_{uniform}+\mathcal{L}_{\phi}$ (8)
Figure 2: An illustration of improved Hourglass architecture. (a) illustrates the overall framework, where we add $c_{flag}$, an additional light-aware classifier to represent $\phi$. The classifier heads serve as initial heads for predicting tokens. (b) illustrates that we change the initial RoPE to a multi-level RoPE, and (c) illustrates the multi-level RoPE.
Improved Hourglass framework. We improve the Hourglass framework via leveraging multi-level information (face-vertex-sequence). Meanwhile. we will show a trick to unify three models $\theta_{mask}$, $\theta_{uniform}$ and $\phi$ into one model $\theta$ with a light-aware classifier head in $\theta$ (Also noted as $\phi$ to keep consistency). The overall framework of the improved Hourglass framework is shown in Fig. 3.
Firstly, multi-level information exists in two places: 1) the Hourglass block, since Hourglass will compress the token sequence into sequence-vertex-face level first, and then reconstruction from face-vertex-sequence again. 2) The token position itself. For the Hourglass block, we add extra cross-attention and use the uncompressed level as the condition. This can help the model leverage the previous level to compress and reconstruct, potentially breaking the information bottleneck.
Then, for the token position, we proposed Multi-level RoPE. The non-AR nature of DDMs allows every token to attend to every other token, creating an opportunity to inject richer structural information than is possible in AR models. Standard positional embeddings, such as 1D RoPE, treat the mesh sequence as a flat list, ignoring its inherent hierarchical structure (coordinates within vertices, vertices within faces). To capitalize on DDM’s full attention scope, we propose a novel multi-level RoPE that explicitly encodes the mesh’s hierarchical structure, shown in Fig. 3 (c). Instead of a single position index, each token is assigned a 3D-like positional tuple: $(\text{face\_index},\text{vertex\_index},\text{coordinate\_index})$. This allows the model’s attention mechanism to learn relationships not just between adjacent tokens in the sequence, but also between vertices within the same face, or across different faces, providing a powerful inductive bias for learning topology.
In the end, we show a trick to unify three models into one with a light-aware classifier to leverage the latent from the transformer. We first added a new condition called the conditional task flag, noted as $c_{flag}$. The flag indicates the current task being performed (e.g., $c_{flag}=0$ for mask-based denoising, $c_{flag}=1$ for uniform-based denoising. This approach promotes parameter efficiency, allowing the model to learn shared representations of mesh structure that benefit both topology sculpting and shape refining tasks. This enables us to use the one model $\theta$ to cover the ability of $\theta_{mask}$ and $\theta_{uniform}$. Then, we integrate a light-aware classifier $\phi$ directly into the architecture. A lightweight classification head is attached to the final layer of the Hourglass transformer. This head takes the model’s final token representations as input and predicts the probability of each token being part of a topologically incorrect region. This shared representation ensures that the features learned for generation are effectively reused for error detection.
Connection loss. After enabling the token parallel generation, we find that topology can be represented by correctly repeating the token among shared vertices. Since each token represents a discrete XYZ token, shared vertices will have the same token. Therefore, we leverage this to propose the connection loss as a hard constraint to improve the accuracy of token prediction further.
Firstly, we define the set of all valid shared vertex groups in a batch $b$ as $\mathcal{G}_{b}=\{G_{b,g}\}_{g=1}^{N_{groups,b}}$, where $G_{b,g}$ contains the starting token indices $s_{k}$ for all occurrences of the $g$-th shared vertex. For a token sequence of length $S$ and vocabulary size $V$, let $P_{b,s}\in\mathbb{R}^{V}$ denote the predicted logits for token $s$ in batch $b$.
The connection loss $\mathcal{L}_{connection}$ encourages a consensus in predicted logits for each coordinate dimension (X, Y, Z) across all tokenized occurrences of the same vertex. For each shared vertex group $G_{b,g}\in\mathcal{G}_{b}$, and for each coordinate dimension $j\in\{0,1,2\}$ (corresponding to X, Y, Z), we extract the predicted logits $L_{b,g,j}\in\mathbb{R}^{|G_{b,g}|\times V}$ for all its occurrences:
$L_{b,g,j}=\{P_{b,(s_{k}+j)}\mid s_{k}\in G_{b,g}\}$ (9)
Then, For each $L_{b,g,j}$, the consensus logits $C_{b,g,j}\in\mathbb{R}^{V}$ are computed as the mean of its members:
$C_{b,g,j}=\frac{1}{|G_{b,g}|}\sum_{s_{k}\in G_{b,g}}P_{b,(s_{k}+j)}$ (10)
The consensus loss for a single group $G_{b,g}$ is the average L1 distance of individual logits from their dimension-wise consensus, averaged across the three dimensions:
$\mathcal{L}_{group}(G_{b,g})=\frac{1}{3}\sum_{j=0}^{2}\frac{1}{|G_{b,g}|}\sum_{s_{k}\in G_{b,g}}\left\|P_{b,(s_{k}+j)}-C_{b,g,j}\right\|_{1}$ (11)
The final $\mathcal{L}_{connection}$ is the average of group-level losses across all valid groups in the batch:
$\mathcal{L}_{connection}=\frac{\sum_{b=1}^{B}\sum_{g=1}^{N_{groups,b}}\mathcal{L}_{group}(G_{b,g})}{\sum_{b=1}^{B}N_{groups,b}}$ (12)
This loss directly penalizes topological errors, such as non-manifold edges or disconnected vertices, by ensuring that all predicted coordinate values for a single vertex converge to a consistent value, thereby bootstrapping topological integrity.
In this way, the final loss could be calculated as:
$\mathcal{L}=\mathcal{L}_{mask}+\mathcal{L}_{uniform}+\mathcal{L}_{\phi}+\mathcal{L}_{connection}.$ (13)
We summarize our hybrid inference and decoupling training via improved Hourglass in Algorithm 1 and Algorithm 2. In this way, TSSR successfully maintains accurate token prediction via a parallel token generation manner. Our design enables TSSR to generate meshes with high resolution and up to 10,000 faces.
Algorithm 1 Hybrid inference
1:Input: $\theta$$\phi$$\sigma$, $T$, $F$
2:Output: $x_{1}$
3:Initialize $x_{0}$ by Eq. 1 via mask token.
[…]
5:for $t$ in $0,...\frac{t}{T},\frac{t+\Delta t}{T},...,1]$ do
6:  Calculate $p_{1|t}(x_{t,mask},t,c,c_{flag}=1,\theta)$.
7:  $x_{1|t,mask}\sim p_{1|t}(x_{t,mask},t,c,c_{flag}=1,\theta)$.
8:  $x_{t,uniform}\leftarrow x_{1|t,mask}$.
9:  Calculate $p_{1|t}(x_{t,uniform},t,c,c_{flag}=0,\theta)$.
10:  $x_{1|t,uniform}\sim p_{1|t}(x_{t,uniform},t,c,c_{flag}=0,\theta)$.
11:  if $t=1$ then
12:   $x_{1}\leftarrow x_{1|t,uniform}$.
13:  else
14:   Calculate $x_{t+\Delta t}$ via re-masking. $\triangleright$ Eq. 4
15:  end if
16:end for
17:Return: $x_{1}$
Algorithm 2 Decoupling training
1:Input: Epochs
2:Output: $\theta$$\phi$
3:Initialize $\theta$ and $\phi$
4:for step in $0,...,\text{Epochs}]$ do
5:  Sample $\hat{x}_{1}$ and $c$ from Dataset.
6:  $t\sim[0,1000]$.
7:  Calculate $x_{t,mask}$ via Eq. 2.
8:  Calculate $p_{1|t}(x_{t,mask},t,c,c_{flag},\theta)$
9:  Calculate $\mathcal{L}_{mask}$ via Eq. 5.
10:  $x_{t,uniform}\sim p_{1|t}(x_{t,mask},t,c,c_{flag},\theta)$
11:  Calculate $p_{1|t}(x_{t,uniform},t,c,c_{flag}$
12:  Calculate $\mathcal{L}_{phi}$ via Eq. 7.
13:  Calculate $L_{connection}$ via Eq. 12.
14:  $\theta,\phi\leftarrow\nabla_{\theta,\phi}\mathcal{L}$
15:end for
16:Return: $\theta$$\phi$
5 Experiment
5.1 Implementation
Figure 3: Qualitative results about the generated mesh. It can be found that the shape of the AR-based method will lose some parts. TSSR can benefit greatly from token parallel generation. Meanwhile, the topology of TSSR maintains a similar quality to that of AR’s methods.
Dataset. The dataset consists of 400K meshes, which were collected from Objaverse-XL [6], ShapeNet [2], and other licensed datasets.
Setting. We train a 2B model from scratch. Training is conducted over 20 days on a cluster of 32 NVIDIA A100 40GB GPUs, initially focusing on meshes with face counts in the range $[0,7000]$ to obtain an initial checkpoint. Then, it scales to $[7000,10000]$ face counts trained on the 7 days using a cluster of 16 NVIDIA H20 96GB GPUs. We use a linear learning rate scheduler with a warm-up strategy and the AdamW optimizer [16], increasing from $0,1\times 10^{-4}]$ over the first 100 steps and maintaining $1\times 10^{-4}$ thereafter. We set $T=200$ during inference.
Evaluation metrics. Following previous works [26], we evaluate the quality and fidelity of the mesh using the Hausdorff Distance (HD), Chamfer Distance (CD), Normal Consistency (NC), and F-score (F1).
Baseline. We directly compare our method with the latest ARs method since ARs currently largely outperforms the diffusion models [26]. The baselines include: MeshAnythingV2, BPT [24], TreeMeshGPT [12], and DeepMesh [27]. We exclude the splitting part-like methods, such as MeshMosaic [26], to ensure a fair comparison since the splitting part can cause better topology [25].
5.2 Experimental Results
Table 1: Quantitative results on ShapeNet, Thingi10K [29], and Objaverse. The best scores are emphasized in bold. The second score is emphasized in lines. * represents the resolution is 512.
Dataset Method HD $\downarrow$ $\mathrm{CD}_{L1}\downarrow$ $\mathrm{CD}_{L2}\left(\times 10^{3}\right)\downarrow$ $\mathrm{NC}\uparrow$ F1 $\uparrow$
ShapeNet MeshAnythingV2* 0.078 0.009 0.640 0.911 0.652
BPT* 0.017 0.003 0.012 0.962 0.875
TreeMeshGPT* 0.161 0.034 5.430 0.841 0.556
DeepMesh* 0.037 0.004 0.060 0.967 0.791
TSSR (Our) 0.050 0.013 0.035 0.954 0.976
Thingi10K MeshAnythingV2* 0.167 0.021 2.492 0.842 0.358
BPT* 0.157 0.035 7.771 0.875 0.496
TreeMeshGPT* 0.233 0.060 18.086 0.788 0.387
DeepMesh* 0.165 0.026 3.331 0.853 0.321
TSSR (Our) 0.106 0.032 1.183 0.969 0.775
Objaverse MeshAnythingV2* 0.118 0.015 1.213 0.859 0.430
BPT* 0.151 0.034 7.016 0.846 0.502
TreeMeshGPT* 0.237 0.057 10.507 0.784 0.308
DeepMesh* 0.111 0.016 1.712 0.866 0.471
TSSR (Our) 0.101 0.014 0.863 0.959 0.744
Table 2: Ablation study of the face counts. S1-S3 represent the different strategies to perturb face count.
Dataset Strategy HD $\downarrow$ $\mathrm{CD}_{L1}\downarrow$ $\mathrm{CD}_{L2}\left(\times 10^{3}\right)\downarrow$ $\mathrm{NC}\uparrow$ F1 $\uparrow$
ShapeNet S1 0.050 0.013 0.035 0.954 0.976
S2 0.054 0.012 0.347 0.841 0.970
S3 0.161 0.034 5.430 0.841 0.959
Thingi10K S1 0.106 0.032 1.183 0.969 0.775
S2 0.124 0.032 1.240 0.915 0.731
S3 0.233 0.060 18.086 0.788 0.611
Objaverse S1 0.101 0.014 0.863 0.959 0.744
S2 0.156 0.034 1.878 0.890 0.731
S3 0.237 0.057 10.507 0.784 0.641
We report the comprehensive quantitative results shown in Table 1. TSSR achieves competitive results as an AR-like method. Concretely, in three datasets, TSSR achieves the best performance on all metrics. To further confirm the quality of the meshes, we present additional qualitative results in Fig. 3. It can be found that TSSR can generate artist-like meshes without the topology errors.
Meanwhile, thanks to parallel token generation, the generated mesh will be more complete than the ARs. Since some meshes generated by ARs will lose some necessary faces. In this way, the experimental results prove the validity of TSSR.
Computation cost. To demonstrate the efficiency of the parallel token generation, we compare the computation cost of TSSR for generating 2k faces meshes with BPT and DeepMesh, shown in Table 3.
Table 3: Ablation study about computation cost on a single H20.
Method Times (s)
DeepMesh 73
BPT 79
TSSR (Our) 51
5.3 Ablation Study
Components on Backbone. To illustrate the influence of each improvement for the improved Hourglass transformer, we report the ablation study shown in Table 4. It can be observed that adding cross-attention and multi-level RoPE can improve the quality of generated text.
Table 4: Ablation study about components in Hourglass on ShapeNet. N/A indicates that the model cannot converge within the same number of training steps.
Cross Attention RoPE HD $\downarrow$ $\mathrm{CD}_{L1}\downarrow$
N/A N/A
$\surd$ 0.1338 0.025
$\surd$ $\surd$ 0.050 0.013
Table 5: Ablation study about connection loss on ShapeNet. W and W/O represent whether the connection loss is used during training, respectively.
Connection Loss HD $\downarrow$ $\mathrm{CD}_{L1}\downarrow$
W 0.122 0.024
W/O 0.050 0.013
Figure 4: Qualitative results about the ablation study of connection loss.
Ablation study about connection loss. To illustrate the importance of connection loss, we present an ablation study in Table 5 and related qualitative findings in Fig. 4. It can demonstrate that the connection can enhance the topology quality. After canceling, the generated mesh will contain some fragile triangles. These triangles are in the correct place according to the cloud point condition, but they generate an invalid mesh.
Face counts. We show a powerful ability of TSSR that can precisely control the face counts in Table 2. We implement three strategies: S1) Using the GT face count. S2) Introducing bias (the face counts will randomly increase/decrease among 100 faces). S3) Maximize bias (the face counts will keep a constant value, i.e., 2k). We also report detailed qualitative results in the supplementary material. Interestingly, TSSR can generate topologies that are entirely different yet artist-like via different face counts, demonstrating its potential.
6 Conclusion
In this paper, we propose TSSR, the DDM-based parallel token generation for direct mesh generation tasks. The motivation of TSSR is to maintain high token prediction accuracy after enabling token parallel generation. TSSR is composed of three novel components: 1) the decouple training and hybrid inference process. 2) An improved Hourglass architecture. 3) A novel connection. Unify them into one, TSSR successfully generates artist-like meshes via parallel token generation. Experimental results demonstrate that TSSR can generate 10K face meshes up to 1024 resolution.
Limitation. Similar to the DDM-based LLM, TSSR also suffers from inconvenient training. For example, we need to pad the sequence length to accommodate the large batch size, which will inevitably incur some computational cost. Such a problem could be resolved via infilling, which remains future work.